OpenStudy (anonymous):

how to factor 16x^4-625

OpenStudy (whpalmer4):

Here's a hint: the difference of two squared quantities can be factored as follows: \[a^2-b^2 = (a-b)(a+b)\]

OpenStudy (whpalmer4):

You may be able to apply that rule more than once in this problem!

OpenStudy (anonymous):

wait never mind I think I got it, would it be (2x^2-5)?

OpenStudy (whpalmer4):

Well, let's make all the pieces match between \[16x^4-625\]and\[a^2-b^2\]We'll set \[a^2=16x^4\]and \[b^2=625\] What are the values of \(a\) and \(b\)?

OpenStudy (anonymous):

a=2 and b=5 right?

OpenStudy (anonymous):

a=4 and b=25

OpenStudy (whpalmer4):

yes on \(b\), no on \(a\)

OpenStudy (anonymous):

4x^2 is a

OpenStudy (whpalmer4):

That's better :-)

OpenStudy (anonymous):

so it would be (4x^2-25) (2x-5) (2x+5) right?

OpenStudy (anonymous):

i mean +25

OpenStudy (whpalmer4):

So, if \(a = 4x^2\) and \(b = 25\), we can rewrite our polynomial as \[16x^4-625 = a^2-b^2 = (a-b)(a+b) = (4x^2-25)(4x^2+25)\]

OpenStudy (whpalmer4):

Yes! You correctly spotted that one of the factors was also a difference of squares, great!

OpenStudy (anonymous):

thank you!

OpenStudy (whpalmer4):

There's also a formula for the difference of cubes, but it isn't quite so neat and tidy.

OpenStudy (whpalmer4):

\[a^3-b^3 = (a-b)(a^2+ab+b^2)\]

OpenStudy (anonymous):

umm idk

OpenStudy (whpalmer4):

@tanzie32 what don't you know?